If you borrowed 10, 000 pesos from a bank with 18% interest per annum, what is the total amount to be repaid at the end of one year?
Answer by ikleyn(52776)

.

Then the total amount to be repaid at the end of one year is  10000 + 0.18*10000 = 10000 + 1800 = 11800 pesos.    ANSWER

Solved.

---------------

To see many other similar  (and different)  solved problems on simple interest accounts,  look into the lesson
    - Simple interest percentage problems
in this site.


Learn the subject from there.